OpenStudy (anonymous):

Find the sum of the series: the problem looks like this, a 4 over the big E looking thing (which I can't show at this time) i = -1 below, 6i - 5 right next to it.

OpenStudy (anonymous):

\[\sum_{}^{4}( 6i-5), i = \sqrt{-1}?\]

OpenStudy (anonymous):

\[\sum_{i=1}^{4}\] 6i - 5

OpenStudy (anonymous):

The 6i - 5 is right next to it not below it

OpenStudy (anonymous):

(6*1-5) +(6*2-5) +(6*3-5) +(6*4-5) = 40

OpenStudy (anonymous):

No wonder! I was looking everywhere on how to solve these type of problems. Thank you, so the answer comes out to be 40 correct?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Thanks again :D

OpenStudy (anonymous):

the "big e" is called Sigma, remember sigma for SUM, have fun

OpenStudy (anonymous):

Sigma, got it